Suppose you have been told that the marginal propensity to consume for Kenya is 0.75. By how much will consumption and savings change if aggregate expenditure (i.e. Disposable income) is increased by ksh.1million?

1
Expert's answer
2021-03-08T09:04:58-0500

Calculating multiplier with MPC the formula for multiplier with MPC is "MPC=1\/(1-MPC)" So multiplier is "1\/(1-0.75)=1.33" or "4\/3" .

The Marginal Propensity to save (MPS) is given as "MPS=(1-MPC)" . So MPS is given as "1-0.75=0.25" . Multiplier with MPS is given as "1\/mps =1\/0.25 = 4\/3 or 1.33" . Change in aggregate demand is "4\/3*1000000 = 1300000" . Change in consumption is "0.75*1000000 = 750,000" .

Change in savings is "0.25 * 1000000 = 250,000" .

Thus consumption and savings changes by ksh750000 and ksh250000 respectively.

Understanding MPS and MPC

For every I ksh consumption changes by 75 cents when MPC is 0.75

For every 1 ksh, 25 cents are saved when MPS is 0.25